OpenStudy (anonymous):

Solve 1/7 + 5/6 Then use estimation to check for reasonableness. (Show your work).

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\rightarrow {1 \over 7} + {5 \over 6} = {6 \over 42} + {35 \over 42} }\)

Parth (parthkohli):

After getting the fractions as like, you must add the numerators.

jhonyy9 (jhonyy9):

in the first step you need getting for 6 and 7 the common denominator so what will be 6*7=42 - than for this case you need multiplie the numerator of 7 by 6 and the numerator of 6 by 7 - so in this way will resulted 1/7 +5/6 = (1*6)/42 + (5*7)/42 =(6+35)/42 = 41/42